# Janus-faces of Temporal Constraint Languages: A dichotomy of expressivity

### v1

@@ 5.3. Min-clean tuples for $\operatorname{mx}$
Let $E$ be a binary relation on $\mathbb{Q}^{k}$. If $E$ is preserved by $\operatorname{mx}$, then either $E$ contains a tuple $t$ with $|M(t)|=1$, or for all $t,t^{\prime}\in E$ with $\min(t_{i})=\min(t^{\prime}_{i})$ and $\operatorname{\textnormal{minx}}(t_{i})=\operatorname{\textnormal{minx}}(t^{\prime}_{i})$ for $i\in\{1,2\}$, also $\operatorname{\textnormal{minx}}(t_{j})=\operatorname{\textnormal{minx}}(t^{\prime}_{j})$ for $j\neq i$.
@@
Suppose that all tuples $t$ satisfy $|M(t)|=2$. Without loss of generality, let $t,t^{\prime}\in E$ be such that $c\coloneqq\min(t_{1})=\min(t^{\prime}_{1})$ and $\operatorname{\textnormal{minx}}(t_{1})=\operatorname{\textnormal{minx}}(t^{\prime}_{1})$. Observe that for all $a,b\in\mathbb{Q}^{k}$ with $\min(a)=\min(b)$ it holds that $\min(\operatorname{mx}(a,b))=\beta(\min(a))$ if and only if $\operatorname{\textnormal{minx}}(a)\triangle\operatorname{\textnormal{minx}}(b)=\varnothing$, where $\beta$ is the endomorphism in the definition of $\operatorname{mx}$. The tuple $s\coloneqq\operatorname{mx}(t,t^{\prime})$ now satisfies $\min(s_{1})=\beta(c)$. By assumption, we have $\min(s_{2})=\min(s_{1})=\beta(c)$, and thus indeed $\operatorname{\textnormal{minx}}(t_{2})\triangle\operatorname{\textnormal{minx}}(t_{2}^{\prime})=\varnothing$. ∎
@@
Let $E$ be a binary relation on $\mathbb{Q}^{k}$. If $E$ has pseudo-algebraic length $1$ and is preserved by $\operatorname{mx}$, then $E$ contains a min-clean tuple.
@@
Any tuple with $|M(t)|=1$ is min-clean. Suppose that $|M(t)|=2$ for all $t\in E$. Take $W$ a weakly connected component of pseudo-algebraic length $1$. As the $\sim_{k}$-factor of the relation $E|_{W\times W}$ is finite, smooth, weakly connected, and of algebraic length $1$, it is $m$-linked for some $m\in\mathbb{N}$. We claim that every $t\in E|_{W\times W}$ is min-clean. Indeed, for all $t\in E|_{W\times W}$ there are $t^{\prime}_{2}\sim_{k}t_{2}$ and an $m$-fence in $E|_{W\times W}$ connecting $t_{1}$ and $t^{\prime}_{2}$. Similar to the proof of Lemma 5.4, repeatedly applying Lemma 5.8 instead of Lemma 5.3, it follows that $\operatorname{\textnormal{minx}}(t_{1})=\operatorname{\textnormal{minx}}(t_{2})$.
@@
Let $R$ be an $n$-ary 22-transitive relation on $\mathbb{Q}^{k}$ with $n\geq 3$. If $R$ is preserved by $\operatorname{mx}$, then $R$ contains a min-clean tuple.
@@
Suppose that $t\in R$ is not min-clean. In particular, we must have $|M(t)|>1$. Without loss of generality, assume $1,2\in M(t)$ and $\operatorname{\textnormal{minx}}(t_{1})\neq\operatorname{\textnormal{minx}}(t_{2})$. Observe that $\operatorname{\textnormal{minx}}(\operatorname{mx}(t_{1},t_{2}))=\operatorname{\textnormal{minx}}(t_{1})\triangle\operatorname{\textnormal{minx}}(t_{2})$, and $\min(\operatorname{mx}(t_{1},t_{2}))=\alpha(\min(t))$. By $2$-transitivity of $R$, there is a tuple $t^{\prime}\in R$ for which $t^{\prime}_{1}=t_{2}$ and $t^{\prime}_{2}=t_{1}$. The tuple $s\coloneqq\operatorname{mx}(t,t^{\prime})$ is contained in $R$, and it satisfies $M(s)=\{1,2\}$, $\min(s)=\alpha(\min t)$, and $\operatorname{\textnormal{minx}}(\textnormal{pr}_{1}(s))=\operatorname{\textnormal{minx}}(\textnormal{pr}_{2}(s))$. ∎
@@ 5.4. Min-clean tuples for $\ell\ell$
Let $R$ be preserved by $\operatorname{\textnormal{lex}}$, $t^{1}\in R$ min-ready, and $t^{2},\dots,t^{m}\in R$. If $i\in M(t^{1})\cap\dots\cap M(t^{m}),$ $\operatorname{\textnormal{minx}}(t^{1}_{i})\cap\dots\cap\operatorname{\textnormal{minx}}(t^{m}_{i})\neq\varnothing$, and $j\in M(t_{1})$, then also $j\in M(t^{2})\cap\dots\cap M(t^{m})$ and $\operatorname{\textnormal{minx}}(t^{1}_{j})\cap\dots\cap\operatorname{\textnormal{minx}}(t^{m}_{j})\neq\varnothing.$
@@
Consider the tuple $t$ defined by $t\coloneqq{\operatorname{\textnormal{lex}}^{[m]}}(t^{1},\dots,t^{m})$. Observe that $\operatorname{\textnormal{minx}}(t_{i})=\operatorname{\textnormal{minx}}(t^{1}_{i})\cap\dots\cap\operatorname{\textnormal{minx}}(t^{m}_{i})$ and $\min(t)={\operatorname{\textnormal{lex}}^{[m]}}(\min(t^{1}_{i}),\dots,\min(t^{m}_{i}))$. By min-readiness of $t^{1}$, we also have $\min(t_{i})=\min(t_{j}),$ which by injectivity of $\operatorname{\textnormal{lex}}^{[m]}$} implies $\operatorname{\textnormal{minx}}(t^{1}_{j})\cap\dots\cap\operatorname{\textnormal{minx}}(t^{m}_{j})\neq\varnothing.$
@@
Let $E$ be a binary relation on $\mathbb{Q}^{k}$. Suppose that $W$ is a weakly connected component of $E$ of pseudo-algebraic length $1$. If $E$ is smooth and preserved by $\operatorname{\textnormal{lex}}$, then either $E$ contains a tuple $t$ with $|M(t)|=1$, or $\bigcap_{a\in W}\operatorname{\textnormal{minx}}(a)\neq\varnothing$.
@@
Let $E$ be a binary relation on $\mathbb{Q}^{k}$ that is smooth and preserved by $\operatorname{\textnormal{lex}}$. If $S\subseteq E$ is smooth itself, has pseudo-algebraic length $1$, and $\textnormal{pr}_{1}(S)$ is weakly connected, then $S^{\prime}$ contains a min-clean tuple.
@@
By Lemma 3.7, $S^{\prime}$ is non-empty. If $E$ contains a tuple $t$ with $|M(t)|=1$, then for every $s\in S^{\prime}$ the tuple $\operatorname{\textnormal{lex}}(t,s)$ is min-clean and contained in $S^{\prime}$.
Suppose now that $|M(t)|=2$ for all $t\in E$. By connectedness, all tuples in $S$ share the same minimal entry $c$. Choose $t^{1},\dots,t^{m}\in S$ such that every $\sim_{2k}$-class of an element in $S$ coincides with the $\sim_{2k}$-class of one of the tuples among $t^{1},\dots,t^{m}$. Once more, by smoothness of $S$, the set of all $\sim_{k}$-classes of the tuples $t^{1}_{1},\dots,t^{m}_{1}$ coincides with the set of all $\sim_{k}$-classes of the tuples $t^{1}_{2},\dots,t^{m}_{2}$. Thus, by the choice of $t^{1},\dots,t^{m}$, we have
$\bigcap_{i\in[m]}\operatorname{\textnormal{minx}}(t_{1}^{i})=\bigcap_{i\in[m]}\operatorname{\textnormal{minx}}(t_{2}^{i})=\bigcap_{a\in\textnormal{pr}_{1}(S)}\operatorname{\textnormal{minx}}(a).$
Lemma 5.12 applied to the weakly connected component of $E$ containing $\textnormal{pr}_{1}(S)$ yields $\bigcap_{a\in\textnormal{pr}_{1}(S)}\operatorname{\textnormal{minx}}(a)\neq\varnothing.$ Therefore, the tuple $s\coloneqq{\operatorname{\textnormal{lex}}^{[m]}}(t^{1},\dots,t^{m})$ satisfies $\min(s_{1})=\min(s_{2})={\operatorname{\textnormal{lex}}^{[m]}}(c,\dots,c)$, as well as $\operatorname{\textnormal{minx}}(s_{1})=\bigcap_{a\in\textnormal{pr}_{1}(S)}\operatorname{\textnormal{minx}}(a)=\operatorname{\textnormal{minx}}(s_{2})$. Moreover, by the choice of $t^{1},\dots,t^{m}$ and injectivity of $\operatorname{\textnormal{lex}}^{[m]}$}, we also have $s\in S^{\prime}$. ∎
@@
Let $R$ be an $n$-ary 22-transitive relation on $\mathbb{Q}^{k}$ that is preserved by $\operatorname{\textnormal{lex}}$. Either there is $t\in R$ with $|M(t)|=1,$ or $|M(t)|=n$ for all $t\in R$.
@@
The proof is similar to the one of Lemma 5.6, using $\operatorname{\textnormal{lex}}^{[n-1]}$} in place of $\operatorname{mi}^{[n-1]}$}. ∎
@@
Let $R$ be an $n$-ary relation on $\mathbb{Q}^{k}$ with $n\geq 3$. If $R$ is cyclic, $2$-transitive, and preserved by $\operatorname{\textnormal{lex}}$, then $R^{\prime}$ contains a min-clean tuple.
@@
If $R$ contains a tuple $t$ with $|M(t)|=1$, this tuple is min-clean. It then follows that for any $s\in R^{\prime}$ (which exists by Lemma 3.7), the tuple $\operatorname{\textnormal{lex}}(t,s)\in R^{\prime}$ is also min-clean. If $R$ does not contain such a tuple, then by Lemma 5.14 all tuples satisfy $|M(t)|=n$ and are, in particular, min-ready. Take $t\in R^{\prime}$ arbitrarily. As in the proof of Lemma 5.7, employing Lemma 5.11 instead of Lemma 5.3, one shows that $\bigcap_{i\leq n}\operatorname{\textnormal{minx}}(t_{i})\neq\varnothing.$ Let $t^{1},\dots,t^{n}$ denote all cyclic permutations of the tuple $t$. We have $t^{i}\in R$ for all $i\leq n$ by cyclicity of $R$. By injectivity of $\operatorname{\textnormal{lex}}^{[n]}$}, the tuple $s\coloneqq{\operatorname{\textnormal{lex}}^{[n]}}(t^{1},\dots,t^{n})$ is contained in $R^{\prime}$. Moreover, observe that $\min(s)={\operatorname{\textnormal{lex}}^{[m]}}(\min(t),\dots,\min(t))$ and $\operatorname{\textnormal{minx}}(s_{i})=\bigcap_{j\leq n}\operatorname{\textnormal{minx}}(t_{j})$ for all $i\leq n$. In particular, $s\in R^{\prime}$ is min-clean. ∎

### v2

@@ 6.3. Min-clean tuples for $\operatorname{mx}$
Let $E$ be a binary relation on $\mathbb{Q}^{k}$ that is preserved by $\langle\operatorname{mx}\rangle$. Either $E$ contains a tuple $t$ with $|M(t)|=1$, or for all $t,t^{\prime}\in E$ with $\min(t_{i})=\min(t^{\prime}_{i})$ and $\operatorname{minx}(t_{i})=\operatorname{minx}(t^{\prime}_{i})$ for $i\in\{1,2\}$, also $\operatorname{minx}(t_{j})=\operatorname{minx}(t^{\prime}_{j})$ for $j\neq i$.
@@
Suppose that all tuples $t$ satisfy $|M(t)|=2$. Without loss of generality, let $t,t^{\prime}\in E$ be such that $c\coloneqq\min(t_{1})=\min(t^{\prime}_{1})$ and $\operatorname{minx}(t_{1})=\operatorname{minx}(t^{\prime}_{1})$. Observe that for all $a,b\in\mathbb{Q}^{k}$ with $\min(a)=\min(b)$ it holds that $\min(\operatorname{mx}(a,b))=\beta(\min(a))$ if and only if $\operatorname{minx}(a)\triangle\operatorname{minx}(b)=\varnothing$, where $\beta$ is the endomorphism in the definition of $\operatorname{mx}$. The tuple $s\coloneqq\operatorname{mx}(t,t^{\prime})$ now satisfies $\min(s_{1})=\beta(c)$. By assumption, we have $\min(s_{2})=\min(s_{1})=\beta(c)$, and thus indeed $\operatorname{minx}(t_{2})\triangle\operatorname{minx}(t_{2}^{\prime})=\varnothing$. ∎
@@
Let $E$ be a binary relation on $\mathbb{Q}^{k}$ that is preserved by $\langle\operatorname{mx}\rangle$. If $E$ is smooth and of pseudo-algebraic length $1$, then $E$ contains a min-clean tuple.
@@
Any tuple with $|M(t)|=1$ is min-clean. Suppose that $|M(t)|=2$ for all $t\in E$. Take $S\subseteq E$ that is smooth, weakly connected, and has pseudo-algebraic length $1$. As the $\sim_{k}$-factor of $S$ is finite, smooth, weakly connected, and of algebraic length $1$, it is $m$-linked for some $m\in\mathbb{N}$. We claim that every $t\in S$ is min-clean. Indeed, for all $t\in S$ there are $t^{\prime}_{2}\sim_{k}t_{2}$ and an $m$-fence in $S$ connecting $t_{1}$ and $t^{\prime}_{2}$. Similar to the proof of Lemma 6.2, repeatedly applying Lemma 6.8 instead of Lemma 6.1 to the relation $\langle{S}\rangle_{\operatorname{mx}}\subseteq E$, it follows that $\operatorname{minx}(t_{1})=\operatorname{minx}(t_{2})$.
@@
Let $n\geq 3$, and let $R$ be an $n$-ary relation on $\mathbb{Q}^{k}$ that is preserved by $\langle\operatorname{mx}\rangle$. If $R$ is cyclic, then $R$ contains a min-clean tuple.
@@
Let $t\in R$, and set $c\coloneqq\min(t)$. We choose $\alpha$ and $\beta$ in the definition of $\operatorname{mx}$ such that $\alpha$ fixes $c$, such that $\beta(c)$ is strictly smaller than the second smallest entry of $t$ (if such an entry exists), and such that $\alpha(\beta(c))=\beta(c)$. Denote the cyclic permutations of $t$ by $t^{1},\dots,t^{n}$, and set $s\coloneqq{\operatorname{mx}^{[n]}}(t^{1},\dots,t^{n})$. We claim that $s\in R$ is min-clean. To see this, for every $j\in[k]$ set
$N(j)\coloneqq|\{i\in[n]:\operatorname{pr}_{j}(t_{i})=c\}|.$
Observe that whenever $N(j)\neq 0$, then for all $i\in[n]$ the value of $\operatorname{pr}_{j}(s_{i})$ is either $c$ or $\beta(c)$, depending only on the parity of $N(j)$. Moreover, all remaining entries of $s$ are greater than $\beta(c)$. It follows that $s$ is indeed min-clean. ∎
@@ 6.4. Min-clean tuples for $\ell\ell$
Remarkably, for the existence of min-clean tuples in relations preserved by $\ell\ell$, we will in fact only make use of the binary operation $\operatorname{lex}$ that is generated by $\ell\ell$.
@@
Let $n\geq 2$, and let $R$ be an $n$-ary relation on $\mathbb{Q}^{k}$ that is preserved by $\operatorname{lex}$. If $t^{1}\in R$ is min-ready, and $t^{2},\dots,t^{m}\in R$ satisfy $i\in M(t^{1})\cap\dots\cap M(t^{m}),$ $\operatorname{minx}(t^{1}_{i})\cap\dots\cap\operatorname{minx}(t^{m}_{i})\neq\varnothing$, and $j\in M(t_{1})$, then also $j\in M(t^{2})\cap\dots\cap M(t^{m})$ and $\operatorname{minx}(t^{1}_{j})\cap\dots\cap\operatorname{minx}(t^{m}_{j})\neq\varnothing.$
@@
Consider the tuple $t$ defined by $t\coloneqq{\operatorname{lex}^{[m]}}(t^{1},\dots,t^{m})$. Observe that $\operatorname{minx}(t_{i})=\operatorname{minx}(t^{1}_{i})\cap\dots\cap\operatorname{minx}(t^{m}_{i})$ and $\min(t)={\operatorname{lex}^{[m]}}(\min(t^{1}_{i}),\dots,\min(t^{m}_{i}))$. By min-readiness of $t^{1}$, we also have $\min(t_{i})=\min(t_{j}),$ which by injectivity of $\operatorname{lex}^{[m]}$} implies $\operatorname{minx}(t^{1}_{j})\cap\dots\cap\operatorname{minx}(t^{m}_{j})\neq\varnothing.$
@@
Let $S$ be a binary relation on $\mathbb{Q}^{k}$. If $S$ is smooth, weakly connected, and of pseudo-algebraic length $1$, then either $\langle{S}\rangle_{\operatorname{lex}}$ contains a tuple $t$ with $|M(t)|=1$, or $\bigcap_{a\in\operatorname{supp}(S)}\operatorname{minx}(a)\neq\varnothing$.
@@
Let $S$ be a binary relation on $\mathbb{Q}^{k}$. If $S$ is smooth, weakly connected, and of pseudo-algebraic length $1$, then $S^{\prime}$ contains a min-clean tuple.
@@
By Lemma 3.7, $S^{\prime}$ is non-empty. If $\langle{S}\rangle_{\operatorname{lex}}$ contains a tuple $t$ with $|M(t)|=1$, then for every $s\in S^{\prime}$ the tuple $\operatorname{lex}(t,s)$ is min-clean and contained in $S^{\prime}$.
Suppose now that $|M(t)|=2$ for all $t\in\langle{S}\rangle_{\operatorname{lex}}$. In particular, every element of $\langle{S}\rangle_{\operatorname{lex}}$ is min-ready. By connectedness, all tuples in $S$ share the same minimal entry $c$. Choose $t^{1},\dots,t^{m}\in S$ such that every $\sim_{2k}$-class of an element in $S$ coincides with the $\sim_{2k}$-class of one of the tuples among $t^{1},\dots,t^{m}$. Once more, by smoothness of $S$, the set of all $\sim_{k}$-classes of the tuples $t^{1}_{1},\dots,t^{m}_{1}$ coincides with the set of all $\sim_{k}$-classes of the tuples $t^{1}_{2},\dots,t^{m}_{2}$. Thus, by the choice of $t^{1},\dots,t^{m}$, we have
$\bigcap_{i\in[m]}\operatorname{minx}(t_{1}^{i})=\bigcap_{i\in[m]}\operatorname{minx}(t_{2}^{i})=\bigcap_{a\in\operatorname{supp}(S)}\operatorname{minx}(a).$
Lemma 6.12 yields $\bigcap_{a\in\operatorname{supp}(S)}\operatorname{minx}(a)\neq\varnothing$. Therefore, the tuple $s\coloneqq{\operatorname{lex}^{[m]}}(t^{1},\dots,t^{m})$ satisfies $\min(s_{1})=\min(s_{2})={\operatorname{lex}^{[m]}}(c,\dots,c)$, as well as $\operatorname{minx}(s_{1})=\bigcap_{a\in\operatorname{supp}(S)}\operatorname{minx}(a)=\operatorname{minx}(s_{2})$. Moreover, by the choice of $t^{1},\dots,t^{m}$ and injectivity of $\operatorname{lex}^{[m]}$}, we also have $s\in S^{\prime}$. ∎
@@
Let $n\geq 2$, and let $R$ be an $n$-ary 22-transitive relation on $\mathbb{Q}^{k}$ that is preserved by $\operatorname{lex}$. Either there is $t\in R$ with $|M(t)|=1,$ or $|M(t)|=n$ for all $t\in R$.
@@
The proof is similar to the one of Lemma 6.4, using $\operatorname{lex}^{[n-1]}$} in place of $\operatorname{mi}^{[n-1]}$}. ∎
@@
Let $n\geq 3$, and let $R$ be an $n$-ary relation on $\mathbb{Q}^{k}$ that is preserved by $\langle\operatorname{lex}\rangle$. If $R$ is cyclic and $2$-transitive, then $R^{\prime}\subseteq R$ contains a min-clean tuple.
@@
If $R^{\prime}$ contains a tuple $t$ with $|M(t)|=1$, this tuple is min-clean. If $R^{\prime}$ (which is non-empty by Lemma 3.7) does not contain such a tuple, then by Lemma 6.14 all tuples in $R^{\prime}$ satisfy $|M(t)|=n$ and are, in particular, min-ready. Take $t\in R^{\prime}$ arbitrarily. As in the proof of Lemma 6.5, employing Lemma 6.11 instead of Lemma 6.1, one shows that $\bigcap_{i\leq n}\operatorname{minx}(t_{i})\neq\varnothing.$ Let $t^{1},\dots,t^{n}$ denote all cyclic permutations of the tuple $t$. We have $t^{i}\in R^{\prime}$ for all $i\leq n$ by cyclicity of $R$ and Lemma 3.7. The tuple $s\coloneqq{\operatorname{lex}^{[n]}}(t^{1},\dots,t^{n})$ is contained in $R^{\prime}$. Moreover, observe that $\min(s)={\operatorname{lex}^{[m]}}(\min(t),\dots,\min(t))$ and $\operatorname{minx}(s_{i})=\bigcap_{j\leq n}\operatorname{minx}(t_{j})$ for all $i\leq n$. In particular, $s\in R^{\prime}$ is min-clean. ∎